Chapter 3 Homework A Saved Help 12.5 points Branson paid $512,300 cash for all of the outstanding common stock of Wolfpack, IC. (1) ENTRY S  (2) ENTRY A, (3) ENTRY I  (4) ENTRY D (5) ENTRY E

D. (1) ENTRY C (2) ENTRY S, (3) ENTRY A (4) ENTRY I (5) ENTRY E
Answer #1

Anecer Page No 0 @ Towinal entry to record the acquisition of the shares of Subsidiary welfare incs - Particular Debit I GoePage. No@ Consolidated worksheet entries Equila method Retained carnêngs = $143000 + 60200 - 20,000 = $183200 Particular I Ge- Page No ③ Excess = 562000 – 343000 = $ 219000 Allocation to soyalty agreement = $146000 Geodusill = 219000 – 146000 : $ = 3